# Qipeng Kuang Ondřej Kuželka Yuanhong Wang Yuyi Wang

### v1

@@ 4 $16uv^{6}+15uv^{8}+6uv^{10}+uv^{12}+15u^{2}v^{4}+4u^{2}v^{6}+6u^{3}v^{2}+u^{4}$ 16 38
Both OEIS sequence A001187 and $\mathsf{WFOMC}(\Psi_{ug}\land connected_{1}(R),n,w,\overline{w})$ count the number of connected undirected graphs on $n$ labeled nodes. By Theorem 3.1, we compute $occ_{n}=[u]f_{n}(u-1;\Psi_{ug},w,\overline{w},R)$ for each positive integer $n$ shown in Table 3. The sequence $occ_{n}$ matches the sequence A001187.
@@
Both OEIS sequence A047864 and $\mathsf{WFOMC}(\Psi_{ug}\land bipartite(R),n,w,\overline{w})$ count the number of bipartite graphs on $n$ labeled nodes. By Theorem 3.2, we compute $b_{n}=f_{n}(-\frac{1}{2};(\Psi_{ug})_{b},w,\overline{w},R)$ for each positive integer $n$ shown in Table 3. The sequence $b_{n}$ matches the sequence A047864.
@@
Both OEIS sequence A000272 and $\mathsf{WFOMC}(\Psi_{ug}\land tree(R),n,w,\overline{w})$ count the number of trees on $n$ labeled nodes. By Theorem 3.3, we compute $tr_{n}=[uv^{2(n-1)}]f_{n}(u-1,v;\Psi_{ug},w,\overline{w},R)$ for each positive integer $n$ shown in Table 3. The sequence $tr_{n}$ matches the sequence A000272.
@@
Both OEIS sequence A001858 and $\mathsf{WFOMC}(\Psi_{ug}\land forest(R),n,w,\overline{w})$ count the number of forests on $n$ labeled nodes. By Theorem 3.4, we compute $fo_{n}=\sum_{i\in[n]}[u^{i}v^{2(n-i)}]f_{n}(u-1,v;\Psi_{ug},w,\overline{w},R)$ for each positive integer $n$ shown in Table 3. The sequence $fo_{n}$ matches the sequence A001858.
@@
Both OEIS sequence A003030 and $\mathsf{WFOMC}(\Psi_{dg}\land SC(R),n,w,\overline{w})$ count the number of strongly connected directed graphs on $n$ labeled nodes. By Theorem 4.1, we compute $sc_{n}=-[u]\bar{g}_{n}(u-1,-2;\Psi_{dg},w,\overline{w},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $sc_{n}$ matches the sequence A003030.
@@
Both OEIS sequence A054946 and $\mathsf{WFOMC}(\Psi_{dg}\land SCT(R),n,w,\overline{w})$ count the number of strongly connected tournaments on $n$ labeled nodes. By Example 4.1, we compute $sct_{n}=-[u]g_{n}(u-1,-2;(\Psi_{dg})_{TN},w,\overline{w},R)$ for each positive integer $n$ shown in Table LABEL:tab:oeis-directed. The sequence $sct_{n}$ matches the sequence A054946.
@@
Both OEIS sequence A003024 and $\mathsf{WFOMC}(\Psi_{dg}\land AC(R),n,w,\overline{w})$ count the number of directed acyclic graphs on $n$ labeled nodes. By Theorem 4.1, we compute $ac_{n}=(-1)^{n}\cdot g_{n}(0,-2;\Psi_{dg},w,\overline{w},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $ac_{n}$ matches the sequence A003024.
@@
Both OEIS sequence A082402 and $\mathsf{WFOMC}(\Psi_{dg}\land AC_{1}(R),n,w,\overline{w})$ count the number of weakly connected directed acyclic graphs on $n$ labeled nodes. By Theorem 5.1, we compute $cac_{n}=(-1)^{n}\cdot[u]g_{n}(u-1,-2;\Psi_{dg},w,\overline{w},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $ac_{n}$ matches the sequence A082402.
@@
Both OEIS sequence A000169 and $\mathsf{WFOMC}(\Psi_{dt}\land AC(R),n,w,\overline{w})$ count the number of directed trees on $n$ labeled nodes, where
@@
By Theorem 4.3 and [MBS23, Proposition 14], we compute $dt_{n}=(-1)^{n}\cdot g_{n}(0,-2;\Psi_{dt},w,\overline{w},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $dt_{n}$ matches the sequence A000169.
@@ 8 Conclusion
In this paper, we propose a novel perspective to investigate the weighted first-order model counting problem with axioms through the lens of graph polynomials. We introduce several polynomials, WCP, NSCP, SSCP, and MWCP, to address the challenge of incorporating axioms into the model counting problem. With these polynomials and their efficient computation algorithms, we recover all existing complexity results on axioms and obtain new results on k-connected-components, bipartite, strongly connected, and spanning subgraph axioms, as well as certain combinations of these axioms. Moreover, we demonstrate a tight connection between the weak connectedness polynomial (resp. strong connectedness polynomials) and Tutte polynomial (resp. directed chromatic polynomials), enabling us to prove that any graph permits an efficient algorithm for computing its Tutte polynomial and directed chromatic polynomials as long as it can be expressed by a fixed $\text{C}^{2}$ sentence with cardinality constraints and unary evidence. Our approach contributes to both first-order model counting and enumerative combinatorics.
A natural direction for future work involves identifying additional domain-liftable axioms with our proposed polynomials. Expanding the applicability of the introduced polynomials to more generalized scenarios, such as the multi-relation version of WCP and SCPs, where the relations are not necessarily nested, presents another avenue for exploration. Additionally, within the field of enumerative combinatorics, using the proposed polynomials to discover intriguing integer sequences, similar to those found in the OEIS, and in [SJT+23], also holds a substantial interest. Finally, from the perspective of complexity theory, our approach offers a unified framework for studying the upper bounds, i.e., the domain-liftability, of the model counting problem with axioms. However, the lower bounds for this problem remain open. It is worth exploring whether there exists a general criterion for determining the hardness of the model counting problem with axioms.

### v2

@@ 4 $16uv^{6}+15uv^{8}+6uv^{10}+uv^{12}+15u^{2}v^{4}+4u^{2}v^{6}+6u^{3}v^{2}+u^{4}$ 16 38
Both OEIS sequence A001187 and $\mathsf{WFOMC}(\Psi_{ug}\land connected_{1}(R),n,\mathds{1},\mathds{1})$ count the number of connected undirected graphs on $n$ labeled nodes. By Theorem 3.6, we compute $occ_{n}=[u]f_{n}(u-1;\Psi_{ug},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $occ_{n}$ matches the sequence A001187.
@@
Both OEIS sequence A047864 and $\mathsf{WFOMC}(\Psi_{ug}\land bipartite(R),n,\mathds{1},\mathds{1})$ count the number of bipartite graphs on $n$ labeled nodes. By Theorem 3.7, we compute $b_{n}=f_{n}(-\frac{1}{2};(\Psi_{ug})_{b},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $b_{n}$ matches the sequence A047864.
@@
Both OEIS sequence A000272 and $\mathsf{WFOMC}(\Psi_{ug}\land tree(R),n,\mathds{1},\mathds{1})$ count the number of trees on $n$ labeled nodes. By Theorem 3.10, we compute $tr_{n}=[uv^{2(n-1)}]f_{n}(u-1,v;\Psi_{ug},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $tr_{n}$ matches the sequence A000272.
@@
Both OEIS sequence A001858 and $\mathsf{WFOMC}(\Psi_{ug}\land forest(R),n,\mathds{1},\mathds{1})$ count the number of forests on $n$ labeled nodes. By Theorem 3.11, we compute $fo_{n}=\sum_{i\in[n]}[u^{i}v^{2(n-i)}]f_{n}(u-1,v;\Psi_{ug},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table 3. The sequence $fo_{n}$ matches the sequence A001858.
@@
Both OEIS sequence A003030 and $\mathsf{WFOMC}(\Psi_{dg}\land SC(R),n,\mathds{1},\mathds{1})$ count the number of strongly connected directed graphs on $n$ labeled nodes. By Theorem 4.7, we compute $sc_{n}=-[u]\bar{g}_{n}(u-1,-2;\Psi_{dg},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $sc_{n}$ matches the sequence A003030.
@@
Both OEIS sequence A054946 and $\mathsf{WFOMC}(\Psi_{dg}\land SCT(R),n,\mathds{1},\mathds{1})$ count the number of strongly connected tournaments on $n$ labeled nodes. By Example 4.8, we compute $sct_{n}=-[u]g_{n}(u-1,-2;(\Psi_{dg})_{TN},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in Table LABEL:tab:oeis-directed. The sequence $sct_{n}$ matches the sequence A054946.
@@
Both OEIS sequence A003024 and $\mathsf{WFOMC}(\Psi_{dg}\land AC(R),n,\mathds{1},\mathds{1})$ count the number of directed acyclic graphs on $n$ labeled nodes. By Theorem 4.9, we compute $ac_{n}=(-1)^{n}\cdot g_{n}(0,-2;\Psi_{dg},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $ac_{n}$ matches the sequence A003024.
@@
Both OEIS sequence A082402 and $\mathsf{WFOMC}(\Psi_{dg}\land AC_{1}(R),n,\mathds{1},\mathds{1})$ count the number of weakly connected directed acyclic graphs on $n$ labeled nodes. By Theorem 5.1, we compute $cac_{n}=(-1)^{n}\cdot[u]g_{n}(u-1,-2;\Psi_{dg},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $ac_{n}$ matches the sequence A082402.
@@
Both OEIS sequence A000169 and $\mathsf{WFOMC}(\Psi_{dt}\land AC(R),n,\mathds{1},\mathds{1})$ count the number of directed trees on $n$ labeled nodes, where
@@
By Theorem 4.10 and [MBS23, Proposition 14], we compute $dt_{n}=(-1)^{n}\cdot g_{n}(0,-2;\Psi_{dt},\mathds{1},\mathds{1},R)$ for each positive integer $n$ shown in LABEL:tab:oeis-directed. The sequence $dt_{n}$ matches the sequence A000169.
@@ 7 Conclusion
In this paper, we propose a novel perspective to investigate the weighted first-order model counting problem with axioms through the lens of graph polynomials. We introduce three polynomials, WCP, NSCP and SSCP, to address the challenge of incorporating axioms into the model counting problem. With these polynomials and their efficient computation algorithms, we recover all existing complexity results on axioms and obtain new results on k-connected-components, bipartite, strongly connected, and spanning subgraph axioms, as well as certain combinations of these axioms. Moreover, we demonstrate a tight connection between the weak connectedness polynomial (resp. strong connectedness polynomials) and the Tutte polynomial (resp. directed chromatic polynomials), enabling us to prove that any graph permits an efficient algorithm for computing its Tutte polynomial and directed chromatic polynomials as long as it can be expressed by a set of ground unary literals, cardinality constraints and a fixed $\text{C}^{2}$ sentence. Our approach contributes to both first-order model counting and enumerative combinatorics.
A natural direction for future work involves identifying additional domain-liftable axioms with our proposed polynomials. Additionally, within the field of enumerative combinatorics, using the proposed polynomials to discover intriguing integer sequences, similar to those found in the OEIS, and in [SJT+23], also holds a substantial interest. Finally, from the perspective of complexity theory, our approach offers a unified framework for studying the upper bounds, i.e., the domain-liftability, of the model counting problem with axioms. However, the lower bounds for this problem remain open. It is worth exploring whether there exists a general criterion for determining the hardness of the model counting problem with axioms.